a bag contains red and blue cards. The probability ir drawing red card is 2/5. what is the probability of drawing a blue card.
Answer by Boreal(15235)

This is dichotomous or two possibilities,and their probabilities add to 1.
If p(red)=2/5, p(blue)=3/5.